If tan 20=p, then tan(160-110)=(tan160-tan110)/(1+tan160tan110)=????

Grade:12

1 Answers

Sandeep Bhamoo
5 years ago
Tan(160-110)= {tan(160)+tan(110)} ÷ {1+tan(110)tan(160)} = tan(50)= tan(30+20)= {tan(30)+tan(20)} ÷ {1-tan(30)tan(20)} ={(1/√3)+p}÷ (1+p/√3) = (1+√3p)÷(√3+p). Please do approve my answer if correct